OpenStudy (brit):

the exterior angles of a triangle have measures of (x+10) degrees, (2x+20) and 3x degrees. what is the measure of the smallest interior angle of the triangle

OpenStudy (anonymous):

The sum of the exterior angles must equal 360 degrees. \[(x+10) + (2x+20) + 3x = 360\]\[6x = 360 - 30\]\[x = {330 \over 6}=55\]Now we can plug x into the equation for each external angle. The external angle with the largest magnitude will correspond to the angle with the smallest internal angle.

OpenStudy (brit):

omg thank you! you are a life saver!!!!!!!!!!!!!!!!